OpenStudy (anonymous):

7f+g=5 solve for f **show work**

OpenStudy (anonymous):

7f-g=5-g 7f/7=(5-g)/7 f=(5-g)/7 i believe...

OpenStudy (anonymous):

oh thanks gorgeous

OpenStudy (anonymous):

ax - c = b, for x

OpenStudy (anonymous):

ax-c+c=b+c ax/a=(b+c)a x=(b+c)a

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

rt - 2n=y, for t

OpenStudy (anonymous):

rt-2n+2n=y+2n rt/r=(y+2n)r t=(y+2n)/r